ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 81 Zharinov, Victor V., Steklov Mathematical Institute, Moscow, Russia On derivations of the Heisenberg algebra It is shown that all derivations of the Heisenberg algebra H are inner. Several interesting subalgebras of the Lie algebra 𝔇(H) of all derivations of H are presented.